OpenStudy (anonymous):

1. 5 1/3 + 4 7/9 9 2/9 10 2/9 10 1/9 9 1/9

OpenStudy (hoblos):

5 1/3 + 4 7/9 = 5 + 1/3 + 4 + 7/9 = 9 + 1/3 + 7/9 = 9 + 3/9 + 7/9 = 9 +10/9 = 10 1/9